(1) Let's suppose that a firm's production technology is represented by the following production function. What...

(1)

Let's suppose that a firm's production technology is represented by the following production function.

q space equals space F left parenthesis K comma space L right parenthesis space equals space K space plus space 3 L

What is the marginal product of labor (MPL)?

a. -3 b. 1 c. 3 d. 1/3

What is the marginal product of capital (MPK)?

a. -1 b. 1/3 c. 3 d. 1

What is the marginal rate of technical substitution (MRTS)?

a. 1/3 b.3 c. -1/3 d. 1
Answer #1

Answer 1

(i)

The correct answer is (c) 3.

Production function is given by :

q = F(K , L) = K + 3L

Formula :

MPL = 3

Hence, the correct answer is (c) 3.

(ii)

The correct answer is (d) 1.

Production function is given by :

q = F(K , L) = K + 3L

Formula :

да МРK 1 дк

Hence, the correct answer is (d) 1.

(iii)

The correct answer is (c) -1/3.

As Function is of the form F(K , L). This means that L is on Vertical axis and K is on horizontal axis.

Thus, MRTS = Slope of isoquant = dL/dK = -(MPK/MPL) = -1/3

So, MRTS = -1/3 (This negative sign means that Isoquant is downward sloping)

Hence, the correct answer is (c) -1/3.